OpenStudy (anonymous):

h my! This next one is a rough!! the yearly production of a 6 foot orange tree is 20 lbs. of oranges. A 16ft. tree produces 94 lbs. let h represent the height of an orange tree and p the number of lbs. of oranges produced. Use the information to answer the question. list the ordered pairs what is the ordered pair for the first tree? p-20=37/5_h-6) OR p-20=-37/5(h-6) OR p=37/5h-122/5 OR p-6=5/37(h-20)

OpenStudy (amistre64):

h,p (6,20) is what they are using for the slope point form so its either the first or second one

OpenStudy (amistre64):

10 by 74 is how the slope changes; 74/10 = 37/5 and its positive id go with the first one

OpenStudy (anonymous):

Thank you!! that was right but now its asking what is the predicted production of an 18 foot tree in pounds?

OpenStudy (amistre64):

that is best suited with another form i think... or at least easier to me :)

OpenStudy (amistre64):

p = 37(18)/5 -37(6)/5 +20 = 108.8 perhaps

OpenStudy (anonymous):

It counted that as wrong. :/

OpenStudy (anonymous):

Nevermind, I forgot to round it, lol.

OpenStudy (amistre64):

well I could have made a mistake :)

OpenStudy (amistre64):

lol

OpenStudy (anonymous):

Now its asking the predicted production of a 20 ft. tree in lbs.

OpenStudy (anonymous):

I hate word problems! I ALWAYS get them wrong!

OpenStudy (amistre64):

h = 20; p = 37(20)/5 - 37(6)/5 +20 = 123.6 .. dont forget to round

OpenStudy (amistre64):

p = 7.4h - 24.4 is the standard looking equation ... maybe

OpenStudy (anonymous):

I wish I could keep you for my other assignments. I'm two weeks behind and don't understand the work. I might very well go crazy, lol!

OpenStudy (amistre64):

i plugged the formula into my "ti-83" and got a table to go by now :)

OpenStudy (anonymous):

What is the predicted height(in feet) of a tree that produces 88 lbs. of oranges?

OpenStudy (amistre64):

15 high = 86.6 16 high = 94 so its inbetween there

OpenStudy (anonymous):

Would I add those and divide by 2?

OpenStudy (anonymous):

90.3?

OpenStudy (anonymous):

It counted 90 wrong. :/

OpenStudy (amistre64):

88p = 15.189... tall

OpenStudy (anonymous):

So I would put that as 16?

OpenStudy (anonymous):

It counted 16 wrong.

OpenStudy (anonymous):

Any ideas???